OpenStudy (kayluh123):

The Sugar Sweet Company will choose from two companies to transport its sugar to market. The first company charges $7096 to rent trucks plus an additional fee of $200.50 for each ton of sugar. The second company charges $7500 to rent trucks plus an additional fee $175.25 of for each ton of sugar.

OpenStudy (kayluh123):

For what amount of sugar does the two companys charge the same?

OpenStudy (kayluh123):

what is the cost when the two companys charge the same?

OpenStudy (anonymous):

7096 + 200.50t = 7500 + 175.25t 200.50t - 175.25t = 7500 - 7096 25.25t = 404 t = 404/25.25 t = 16 The two companies charge the same at 16 tons of sugar :)